40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 6 23:29:47 2009, Jenne, Update, PSL, Ref cav reflection PD is funky 
    Reply  Fri Aug 7 00:10:14 2009, rana, Update, PSL, Ref cav reflection PD is funky 
       Reply  Fri Aug 7 17:05:34 2009, Jenne, Update, PSL, Ref cav reflection PD is funky 
Message ID: 1850     Entry time: Thu Aug 6 23:29:47 2009     Reply to this: 1851
Author: Jenne 
Type: Update 
Category: PSL 
Subject: Ref cav reflection PD is funky 

After Alberto and I worked on aligning the reference cavity, Rob asked the important and useful question: what is the visibility of the reference cavity.  This helps tell us if we're optimally aligned or not even close.

I did a scan of the ref cav temperature, using /scripts/PSL/FSS/SLOWscan, but there seems to be no real signal is C1:PSL-FSS_RFPDDC.  As shown in Alberto's 200-day plot, it does change sometimes, but if you zoom in on the flat parts, it seems like it's not really reading anything meaningful.  I did a cursory check-out of it, but I'm not 100% sure where to go from here:  There are (as with all of these gold-box PDs) 3 outputs:  a ribbon cable (for ADC purposes I think), an SMA for the RF signal, and a BNC for the DC signal.  The photodiode is clearly working, since if you stick the Lollypop in front of the PD, the cavity unlocks.  I plugged a 'scope into the DC BNC, and it also behaves as expected: block the beam and the signal goes down; unblock the beam and the signal goes up.  Something of note is that this readout gives a positive voltage, which decreases when the beam is blocked.  However, looking at the dataviewer channel, nothing at all seems to happen when the beam is blocked/unblocked.  So the problem lies somewhere in the get-signal-to-DAQ path.  I unplugged and replugged in the ribbon cable, and the value at which the channel has been stuck changed.  Many days ago, the value was -0.5, for the last few days it's been -1.5, and after my unplug/replug, it's now back to ~ -0.5 . The other day Alberto mentioned, and made the point again today that it's a little weird that the PD reads out a negative voltage.  Hmm.

 

Do we have a tester-cable, so that instead of the ribbon cable, I can plug that connector (or pins thereof) into a 'scope?

ELOG V3.1.3-